OpenStudy (ell):

Solve each equation by completing the square: x^2+12x+20=0

OpenStudy (anonymous):

video :)

OpenStudy (jim766):

start by getting the 20 out of there , subtract it on both sides

OpenStudy (ell):

Thanks @timo86m this is really helping(: @Jim766 so far i got that out of the video and now i have x^2+12 = -20

OpenStudy (jim766):

take the 12 from 12 x and div it by 2 and then square it

OpenStudy (jim766):

12/2 = 6, 6^2 = ?

OpenStudy (ell):

i got 36

OpenStudy (anonymous):

THanks for medal. THe vid is great btw. Tells you how to do it and why it works

OpenStudy (jim766):

right now, add that to both sides x^+12x + 36 = -20 + 36 now simplify

OpenStudy (jim766):

-20 + 36 = ?

OpenStudy (ell):

Np @timo86m thanx for sharing it, it really helped! @Jim766 soo wat u have to do to one side, u do to the other...it would be 16 so x^2+12x+36= 16

OpenStudy (jim766):

right x^2 + 12x + 36 = 16 now the reason we chose 12/2 and then square it is because it makes it very easy to factor x^2 + 12x + 36 = 16 (x + 6)(x + 6) = 16 (x+ 6)^2 = 16 do you understand what I did there?

OpenStudy (ell):

6 is there b/c it goes into both 12 and 36. right?

OpenStudy (jim766):

right 6^2 = 36 and 6+6 = 12

OpenStudy (jim766):

now the reason we did all this was now we have something squared on one side (x + 6)^2 = 36 now take the square root of each side the square root of something squared is the something before it was squared square root of 36 = 6 or -6 so we have x + 6 = + or - 6 you with me on this?

OpenStudy (jim766):

now to finish it we solve both equations x + 6 = 6 or x + 6 = -6 solve both of those for x

OpenStudy (ell):



OpenStudy (jim766):

we have already taken the square root of 36 and got 6, don't need to do it again x + 6 = 6 -6 -6 x = 0 now do the other one

OpenStudy (ell):

So it would be: x + 6 + 6 -6 -6 x=0 x + 6 = -6 -6 -6 x= -12 (0, -12)

OpenStudy (jim766):

sorry I screwed this up it should have been (x+6)^2 = 16 , I put 36 by mistake above so it is x +6 = 4 or x + 6 = -6 -6 -6 -6 -6 x = -2 or x = - 12 sorry for the error, hope I didn't confuse you.